When substituting for a variable in a function graph, how do I select the correct substitution for the variable.
Example: f(x)=-7(x)
When making a graph, I am not sure how to get the input, and output.

Answer by stanbon(75887)

When substituting for a variable in a function graph, how do I select the correct substitution for the variable.
Example: f(x)=-7(x)
------------------------
You give a value to x and find a value for f(x) based on the equation.
For example:
If x = 0, y = -7*0 = 0
If x = 2, y = -7*2 = -14
If x = 100, y = -7*100=-700
======================
Cheers,
Stan H.